core: 0 ol 1 pt roblem 4-3 (similar to) (Evaluating proftabality) L Queston Help | L ast year, Slevens Inc had sales of $402,000, with a cost of goods sold of $111,000. Thesoperating expenses woe $127,900, and is inorease in rotained earnings was 550,000. There are currendty 22,400 comnon słock shares outstandeg and the im pays a $1 56 dividend per share Assuming the firms earnings are taxed at 34 percent consitruct the fems income statement b. Compute the fims operating peofit margin c. What was the times interest eamed? he frmss eamings are taend at 34 percent comstruct tho ms ncome statemert (Rendhstar NOTE You may input eopense accounts as negelive value) Stevens Inc. lncome Statomen 02002 Cost of goods sole1000 291000 127000 Gross profits Operating income (EEIT Earnings before taxes Income taxe Choose from any ist or enter any number n the input hlds and then cck Ceck A cear A Check Answe mainig
Answer #1

1.

Net Income = Increase in Retained Earnings + Dividends paid

= $50,000 + $1.56*22,400

= $84,944

Earnings before Tax = 84,944/(1-0.34) = $128,703

Interest Expense = $164,000 - $128,703

= $35,297

Income Statement will be as follows:

Sales = 402,000

Less: COGS 111,000

Gross Profits 291,000

Operating Expenses 127,000

EBIT 164,000

Less: Interest 35,297

Earnings before taxes = 128,703

Less: Tax expense 43,759

Net Income = $84,944

b.Operating Profit Margin = Operating Profit/Sales

= 164,000/402,000

= 40.8%

c.Times Interest Earned = EBIT/Interest

= 164,000/35,297

= 4.65 times